40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 5 11:49:49 2009, Stephanie, Update, General, Multiply Resonant EOM Update OSASetup.pngSMAPic.pngWithSMA.png
    Reply  Wed Aug 5 15:18:12 2009, Stephanie, Update, General, Multiply Resonant EOM Update SMA.png
Message ID: 1835     Entry time: Wed Aug 5 15:18:12 2009     In reply to: 1834
Author: Stephanie 
Type: Update 
Category: General 
Subject: Multiply Resonant EOM Update 


I have spent the past couple of days gathering optics and mounts so that I can observe the modulation of the EOM attached to the circuit I built using the optical spectrum analyzer (OSA). A rough diagram of the planned layout is attached.

I also built a short SMA cable so that the EOM did not have to be connected directly to the circuit box. The cable is shown attached to the EOM and circuit box in the attached photo. After checking to make sure that all of the connections in the cable were sound, I remeasured the input impedance of the circuit; the impedance measurement (black) is shown in the attached plot with the impedance before the SMA cable was added with and without the box (green and blue, respectively--these two are almost identical). The new impedance has a strange shape compared to the original measurements; I'd like to understand this a little better, since adding extra inductance in LTSpice doesn't seem to have that effect. Since I had already taken apart the setup used for the previous impedance measurements, I had to rebuild and recalibrate the setup; I guess the difference could be something about the new calibration, but I don't really think that that's the case.


After investigating this a bit further, I discovered that some of the components in the circuit were pressed firmly up against the inside of the box, and when they were moved, the impedance plot changed shape dramatically. I think that originally, the components were not pressed against the box, but the box's SMA joint was rather loose; when I connected this to the SMA cable, I tightened it, and this seems to have twisted the circuit around inside the box, pushing the components up against the side. I have fixed the twisting, and since the SMA joint is now tight, the circuit should no longer have any twisting problems.

A new plot is attached, showing the impedance of the circuit with nothing attached (blue), with the SMA cable and EOM attached (green), and with the EOM attached directly to it taken last friday with the old calibration of the setup (red). All three curves look roughly the same; the center peak is shifted slightly between the three curves, but the circuit with SMA and EOM is the version we'll be using, and it's central peak is close to the correct value.

Attachment 1: SMA.png  22 kB  | Hide | Hide all
ELOG V3.1.3-